Find the area between f(x)=x²+4x and g(x)=x²-6x+5 over the interval 0 less or equal to x or x is less or equal to 1


1
Expert's answer
2022-05-09T16:23:22-0400
"x^2+4x=x^2-6x+5"

"x=0.5"

"A=\displaystyle\int_{0}^{0.5}(x^2-6x+5-(x^2+4x))dx"

"+\displaystyle\int_{0.5}^{1}(x^2+4x-(x^2-6x+5))dx"

"=[-5x^2+5x]\begin{matrix}
 0.5 \\
 0
\end{matrix}+[5x^2-5x]\begin{matrix}
 1 \\
 0.5
\end{matrix}"

"=-1.25+2.5+5-5-1.25+2.5"

"=2.5({units}^2)"

Area is 2.5 square units.